# Lodge: A Coarse to Fine Diffusion Network for Long Dance Generation Guided by the Characteristic Dance Primitives

### v2

@@ 1 Introduction
However, generating long dance sequences poses a notable challenge due to the substantial computational resources needed for training. Therefore, many methods are based on autoregressive models [25, 16, 43], and continuously generate dance movements based on a relatively small sliding window. This autoregressive nature accumulates the model prediction errors as time progresses, and prevents the model from learning global choreographic patterns. As a result, motion freezing often occurs after several seconds [57]. There are also some methods [43, 65, 61] maintain a latent space to represent motion, and combine a autoregressive based sequence model to learn music-dance paired relationship. However, the compressed latent space with limited representational capacity also makes these methods prone to overfitting, resulting in poor generalization and diversity. Recently, EDGE [50] proposed a diffusion-based dance generation model. During the denoising process, EDGE parallelly generate multiple dance segments with overlap while maintaining consistency between these overlapping parts using diffusion inpainting  [29], and finally splices these segments into a long dance by linear interpolation. However, their dances lack an overall choreographic structure and shows incoherence at the splices points.
@@ 2.2 Music Driven Dance Generation
In recent years, with the rapid development of neural networks [41, 27, 56, 7, 61], Diffusion makes significant strides in tasks such as image, video, and motion generation [14, 32, 30, 31, 11, 8, 10, 9, 46, 55, 19]. FineDance[26] and EDGE[50] introduce Diffusion to generate diverse and high-quality dance clips of seconds, but they only focus on detailed dance clips and cannot quickly generate long-term dance movements that conform to the overall choreography rules.
@@ 3.1 Preliminaries
Figure 2: An overview of our framework. “TE” is Transformer Encoder, “G” is the genre of dance, “LD” is the Local Diffusion Model.
@@ 3.3 Global Diffusion
Previous work overlooks the global dependencies between music and dance, focusing only on learning the music-dance relationship within a small window. This mainly because long music and dance sequences introduce significant computational overhead. To address this issue, we only task global diffusion to generate sparse dance primitives. Subsequently, multiple local diffusions work in parallel to generate complete long dances.
Given a global music features ${\bm{m}}_{g}$ extracted by the Librosa[34]. We first feed ${\bm{m}}_{g}$ into a two-layer Transformer downsample network, which comprises a Linear layer and a Transformer Encoder Layer. Next, the compressed global music feature is sent to global diffusion. We adopte the EDGE framework as the foundation for global diffusion, making a single modification by adjusting the training objective to output sparse dance primitives. These primitives are key motions with only 8 frames, categorized as ${\bm{d}}_{h}$ and ${\bm{d}}_{s}$.
@@ 4.2 Comparisons on the FineDance dataset
[…]
Motion Diversity. To evaluate the motion diversity of generated dance sequences, we calculate the mean Euclidean distance within the motion feature space, as outlined in the works of Bailando[43]. $\mathrm{DIV}_{k}$ represents the motion diversity in the kinematic feature space, while $\mathrm{DIV}_{g}$ denotes the diversity in the geometric feature space. Table 1 reveals that our Lodge approach achieved the highest $\mathrm{DIV}_{g}$ score, which can be credited to our adoption of global diffusion and characteristic dance primitives for mastering diverse choreography patterns. Concurrently, Lodge’s $\mathrm{DIV}_{k}$ score stands at 6.75, not quite reaching the levels of EDGE and Bailando. This is partly due to our integration of characteristic dance primitives and a foot refine block, which has constrained the solution space for the generated dance.
Beat Alignment Score (BAS). One of the most crucial rules in choreography is the alignment of the dance beat with the music. To evaluate the beat consistency between the generated dance and the given music, we follow [25] and use the BAS to evaluate our methods, our approach demonstrated the highest Beat Alignment Score of 0.2397.
@@
Run Time in Table 1 presents the average Run Time required to generate 1024 frames of dance movements. Bailando achieved the fastest performance, but its runtime increases linearly with the length of the sequence generated. EDGE, using the ddim accelerated sampling strategy and linear interpolation, also achieved a fast level for generating long dance sequences. Our method uses DDPM sampling with a denoising step of 1000, taking 30.93 seconds. Using DDIM with 100 denoising steps takes only 4.57 seconds. Meanwhile, our parallel architecture ensures runtime remains stable even with longer sequences.
[…]
@@ 5 Conclusion and Limitation
In this work, we introduce Lodge, a two-stage coarse-to-fine diffusion network, and propose characteristic dance primitives as intermediate-level representations for the two diffusion models. Lodge has been extensively evaluated through user studies and standard metrics, achieving state-of-the-art results. Our generated samples demonstrate that Lodge can parallelly generate dances that conform to choreographic rules while preserving local details and physical realism. However, our method currently cannot generate dance movements with hand gestures or facial expressions, which are also crucial for performances. We are excited to see the possibility of long-duration full-body dance generation in the future.

### v3

@@ 1 Introduction
However, generating long dance sequences poses a notable challenge due to the substantial computational resources needed for training. Therefore, many methods are based on autoregressive models [22, 14, 40], and continuously generate dance movements based on a relatively small sliding window. This autoregressive nature accumulates the model prediction errors as time progresses, and prevents the model from learning global choreographic patterns. As a result, motion freezing often occurs after several seconds [54]. There are also some methods [40, 62, 58] maintain a latent space to represent motion, and combine a autoregressive based sequence model to learn music-dance paired relationship. However, the compressed latent space with limited representational capacity also makes these methods prone to overfitting, resulting in poor generalization and diversity. Recently, EDGE [47] proposed a diffusion-based dance generation model. During the denoising process, EDGE parallelly generate multiple dance segments with overlap while maintaining consistency between these overlapping parts using diffusion inpainting  [26], and finally splices these segments into a long dance by linear interpolation. However, their dances lack an overall choreographic structure and shows incoherence at the splices frames.
@@ 2.2 Music Driven Dance Generation
In recent years, with the rapid development of neural networks [38, 24, 53, 7, 58], Diffusion-based methods make significant strides in tasks such as image, video, and motion generation [12, 29, 27, 28, 9, 8, 43, 52]. FineDance[23] and EDGE[47] introduce Diffusion to generate diverse and high-quality dance clips of seconds, but they only focus on local motion quality of detailed dance clips and cannot quickly generate long-term dance movements that conform to the overall choreography rules.
@@ 3.1 Preliminaries
Figure 2: An overview of our framework. “TE" is Transformer Encoder, “G" is the genre of dance, “LD" is the Local Diffusion Model.
@@ 3.3 Global Diffusion
Previous works overlook the global dependencies between music and dance, focusing only on the music-dance relationship within a small window. To address this issue, we only task global diffusion to generate sparse dance primitives. Subsequently, multiple local diffusions work in parallel to generate complete long dances.
Given a global music features ${\bm{m}}_{g}$ extracted by the Librosa[31]. We feed ${\bm{m}}_{g}$ into a Transformer downsample network, which comprises a Linear layer and a Transformer Encoder Layer. Next, the compressed global music feature is sent to global diffusion. We adopte the EDGE framework as the foundation for global diffusion, making a single modification by adjusting the training objective to output sparse dance primitives. These primitives are key motions with only 8 frames, categorized as ${\bm{d}}_{h}$ and ${\bm{d}}_{s}$.
@@ 4.2 Comparisons on the FineDance dataset
Motion Diversity. To evaluate the motion diversity of generated dance sequences, we calculate the mean Euclidean distance within the motion feature space, as outlined in the works of Bailando[40]. $\mathrm{DIV}_{k}$ represents the motion diversity in the kinematic feature space, while $\mathrm{DIV}_{g}$ denotes the diversity in the geometric feature space. Table 1 reveals that our Lodge approach achieved the highest $\mathrm{DIV}_{g}$ score, which can be credited to our adoption of global diffusion and characteristic dance primitives for mastering diverse choreography patterns.
Beat Alignment Score (BAS). To evaluate the beat consistency between the generated dance and the given music, we follow [22] and use the BAS to evaluate our methods, our approach demonstrated the highest Beat Alignment Score of 0.2397.
@@
Run Time in Table 1 presents the average Run Time required to generate 1024 frames of dance movements. Bailando achieved a outstanding performance, but its runtime increases linearly with the length of the sequence generated. EDGE, using the DDIM accelerated sampling strategy and linear interpolation, also achieved a fast level for generating long dance sequences. Our method uses DDPM sampling with a denoising step of 1000, taking 30.93 seconds. Using DDIM with 50 denoising steps takes only 4.57 seconds. Meanwhile, our parallel architecture ensures runtime remains stable even with longer sequences.
User study. We conducted a user study where 20 participants viewed 17 video pairs. Each pair consists of two dance sequences: one created by Lodge (DDPM) and the other by different methods or ground truth.
@@ 5 Conclusion and Limitation
In this work, we introduce Lodge, a two-stage coarse-to-fine diffusion network, and propose characteristic dance primitives as intermediate-level representations for the two diffusion models. Lodge has been extensively evaluated through user studies and standard metrics. Our generated samples demonstrate that Lodge can parallelly generate dances that conform to choreographic rules while preserving local details and physical realism. However, our method currently cannot generate dance movements with hand gestures or facial expressions, which are also crucial for performances. This limitation opens avenues for future research.